# Towards a Unified Copernicus Foundation Model for Earth Vision

### v2

@@ Appendix B Copernicus-FM
B.1 Spectral hypernetwork
@@ B.2 Variable hypernetwork
B.3 Metadata integration
@@
Figs. 16 and 17 illustrate the Fourier encoded metadata (location, area, and time) for a few representative example values as below:
  • location ($\text{lon}+180\degree$): $0\degree,45\degree,90\degree,135\degree,180\degree,225\degree,270\degree,315\degree,360\degree,360\degree$;
  • location ($\text{lat}+90\degree$): $0\degree,45\degree,90\degree,135\degree,180\degree,0\degree,45\degree,90\degree,135\degree,180\degree$;
@@
Moving further, we wonder how much benefit each metadata component brings to the model, as well as how the format of each metadata will affect the performance. To answer these questions, Tab. 10 conducts additional ablation on specific metadata components. Results show that geolocation gives the most significant improvement, followed by area and time. Interestingly, geographic coordinates perform better than Cartesian coordinates despite their distortion in high-latitude regions. Using the area corresponding to the true surface coverage (e.g., after cropping and resizing the true surface coverage is smaller) is necessary, without which the performance begins to drop. Using the day of the year and absolute days above one-year-period perform similarly. We use the latter such that it’s convenient to extend to long time series in the future.
@@ B.4 Pretraining details
We pretrain Copernicus-FM on the joint 220K-grid subset of Copernicus-Pretrain, with each grid being one sample unit containing aligned images from all eight modalities. For fast data loading, we convert the raw dataset into webdataset [^1] https://github.com/webdataset/webdataset format, with one grid cell being one minimum sample in the shards. During training, one image from each modality is sampled from one grid cell to construct the input for each iteration. For S1/2, we normalize the image values with channel-wise mean and standard deviation. For S3, we multiply each channel with its corresponding scale factor [^2] https://developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S3_OLCI. For S5P, we use the raw values, and replace NaN pixels with zero. For DEM, we divide the pixel values by 10000. We apply simple data augmentations to each modality, including random resized cropping with scale $[0.2,1.0]$ to its corresponding input size and random horizontal flipping. Each image comes with its metadata, including geolocation (central coordinates in lon/lat), patch area (calculated from GSD and patch size in km2), and time (number of days since a reference date 1970-01-01). The geolocation and patch area are adapted dynamically based on the cropping parameters in data augmentation. Note that despite this adaptation, due to geographical projection the patch area doesn’t strictly reflect the surface area, but is accurate enough for our pretraining purpose. While S1/2/3 images have exact acquisition dates, S5P images are monthly mean and DEM doesn’t have a specific acquisition date. Therefore, we use the first day of the month for one S5P image, and the first day of the year 2015 for all DEM images.
@@ C.2 Benchmark curation
  • Cloud-S3: This is a cloud segmentation dataset with raw images from Sentinel-3 OLCI and labels from the IdePix [66] classification algorithm. We first download a few large cloudy S3 tiles (about $4800\times 400$ pixels) distributed across the globe, and then apply the IdePix algorithm using the ESA SNAP toolbox to get multi-class cloud masks. After that, we manually check the quality of the generated masks, filter out low-quality tiles, and get seven big tiles with high-quality cloud labels. Next, we remap the label IDs, georeference the tiles to GeoTIFFs, and use GDAL to crop the large tiles into small patches with size $256\times 256$ pixels. We remove boundary patches filled with NaN pixels, and split all high-quality patches into 1197/399/399 train/val/test subsets. The class names for the cloud masks are: invalid, clear, cloud-sure, cloud-ambiguous, cloud-shadow, and snow-ice, of which “invalid” should be ignored during training. Apart from the multi-class labels, for each image we also have one binary cloud mask. Therefore, the Cloud-S3 dataset can serve as both a multi-class and a binary segmentation dataset.
@@
  • LC100Cls-S3 and LC100Seg-S3: These two datasets are based on Sentinel-3 OLCI images and CGLS-LC100 [13] land cover maps. CGLS-LC100 is a product in the Copernicus Global Land Service (CGLS) portfolio and delivers a global 23-class land cover map at 100 m spatial resolution. We pick the map product for 2019, and sample and download S3 images and LC100 labels for about 10K locations across the globe using GEE. For each location, we download a land cover map with about $288\times 288$ pixels, and four seasonal S3 OLCI images each with about $96\times 96$ pixels (300 m resolution). Despite using bright pixel percentage to simulate cloud filtering, the resulting images still contain a large volume of clouds. To tackle this issue, we train a cloud detection model based on the previously introduced Cloud-S3 dataset and filter out model-detected cloudy images. After a final quality check, we get about 8K samples each with a land cover map and a time series of S3 images. We divide them into 5181/1727/1727 train/val/test subsets to construct the LC100Seg-S3 dataset. For LC100Cls-S3, we integrate multi-label annotations from the land cover maps for each sample, constructing a multilabel classification dataset. Note that the number of S3 time stamps for different samples may differ because of the cloud filtering process. Apart from the time series, we also pre-select one image for each sample, constructing the “static” version of LC100Cls-S3 and LC100Seg-S3, which is the default mode in Copernicus-Bench.
@@
  • Biomass-S3: This regression dataset is based on Sentinel-3 OLCI images and CCI biomass [52]. The biomass product is part of the European Space Agency’s Climate Change Initiative (CCI) program and delivers global forest above-ground biomass at 100 m spatial resolution. We pick the product for 2020, and the layer of above ground biomass (AGB, unit: tons/ha, i.e. Mg/ha) as regression ground truth, which is defined as the mass, expressed as oven-dry weight of the woody parts (stem, bark, branches and twigs) of all living trees excluding stump and roots. We sample representative regions across the globe, and download corresponding S3 images (one for each season) from GEE and biomass maps from the CCI open data portal. We crop the S3 images into patches with about $96\times 96$ pixels (300 m resolution), and the corresponding biomass maps into patches with about $288\times 288$ pixels. Similar to LC100Cls-S3 and LC100Seg-S3, the resulting S3 images contain a large volume of clouds, thus we use again the cloud detection model to filter out cloudy images. After a final quality check, we acquire 5K samples each with a biomass map and a time series of S3 images. We divide them into 3000/1000/1000 train/val/test subsets to construct the Biomass-S3 dataset. Note that the number of S3 time stamps for different samples may differ because of the cloud filtering process. Apart from the time series, we also pre-select one image for each sample, constructing the “static” version of Biomass-S3, which is also the default mode in Copernicus-Bench.
@@
  • AQ-NO2-S5P and AQ-O3-S5P: These two regression datasets are based on Sentinel-5P NO2 and O3 images and EEA air quality data products [33]. The European Environment Agency (EEA) air quality product provides values for the human health related indicators of air pollutants at 1 km2 grid covering the whole Europe, combining monitoring air quality data in a “regression-interpolation-merging-mapping” methodology and the observational values of the air quality monitoring stations used in the interpolation. We pick the products in 2021 for NO2 (annual average concentration) and Ozone (O3, 93.2 percentile of maximum daily 8-hour means, SOMO35) as regression ground truth. We sample and download S5P NO2 (“tropospheric_NO2_column_number_density”) and O3 (“O3_column_number_density”) images from GEE, and EEA NO2 and O3 maps from EEA datahub [^3] https://www.eea.europa.eu/en/datahub. We use a sample patch size of about $56\times 56$ pixels for both S5P and EEA. For S5P, we download two versions: 1) annual mean, and 2) seasonal mean for each season. After filtering out NaN patches and a final quality check, we get 1480/493/494 train/val/test samples for both NO2 and O3, each with an “annual” mode of 1 S5P image and a “seasonal” mode of 4 S5P images. “Annual” is the default mode in Copernicus-Bench.
@@ C.3 Benchmark characteristics
Figs. 18, 19, 20, 21, 22, 23, 24, 25, 26 and 27 visualize some examples for each dataset in Copernicus-Bench.
@@ C.4 Benchmark implementation
We use simplified data augmentations for training sets: horizontal and vertical flipping for classification tasks, and 90°$\degree$-rotation, horizontal and vertical flipping for segmentation and regression tasks. No augmentation is used for validation and testing sets. Data normalization is performed on the input according to the pretrained model’s preference. For most cases, normalization is performed by subtracting the channel-wise mean and dividing by standard deviation based on the pretrained-model-preferred statistics. If there is no preference, we recommend using the statistics calculated from the training set of each dataset as a standard. For regression tasks, we do mean/std (of the training set) normalization also on the targets to stabilize training. The predicted output is later converted back to the original scale to compute evaluation metrics.
@@ D.1 Climate prediction visualization
Figs. 29 and 31 visualize the prediction results on 10-year mean/std of the six climate parameters, comparing using the geocoordinates or a combination of coordinates and Copernicus-FM-generated grid embeddings as input data. Figs. 30 and 32 further plot the prediction error (Target-Prediction) of using only coordinates, coordinates and embeddings, and only embeddings. The figures show that using raw coordinates captures the general distribution of the climate parameters but tends to be over-smooth, while introducing EO-generated grid embeddings can capture finer details and extremes.

### v3

@@ Appendix B Copernicus-FM
B.1 Subtractive ablation
The incremental ablation in the main paper demonstrates the design evolution, but does not isolate individual contributions. To complement, we conduct a subtractive ablation in Table 8, showing the benefits of each component regardless of order.
Table 8: Subtractive ablation. w/o means without.
EU-S1 EU-S2 EU-RGB LC-S3 O3-S5P ($\downarrow$) Copernicus-FM 81.0 89.5 78.9 90.7 811.6 w/o var hypernet 78.9 87.9 78.6 90.5 857.6 w/o metadata 56.9 88.3 70.1 86.9 1556.3 w/o distill 77.9 88.9 78.5 90.7 839.3
B.2 Spectral hypernetwork
@@ B.3 Variable hypernetwork
Ablation on different variable encoding options
Language encoding maximizes the flexibility to process any variable names, and also maintains semantic relationships between variables. Besides that, random hashing and spectral-sensitivity-guided spectroscopy are two other options to encode different variables. However, compared to language encodings, random hashing loses the flexibility and semantic relationship, while spectroscopy has semantics but lacks flexibility. Quantitatively, we conduct a comparison study in Table 10, suggesting the superior and stable performance of language encoding.
Table 10: Variable name encoding ablation.
EU-S1 EU-S2 EU-RGB LC-S3 O3-S5P ($\downarrow$) LLM (LLaMa-3.2-1B) 81.0 89.5 78.9 90.7 811.6 Random hash 77.7 89.5 78.6 89.9 818.6 Spectroscopy 80.4 89.5 79.5 89.9 815.7
B.4 Metadata integration
@@
Figs. 17 and 18 illustrate the Fourier encoded metadata (location, area, and time) for a few representative example values as below:
[…]
@@
Moving further, we wonder how much benefit each metadata component brings to the model, as well as how the format of each metadata will affect the performance. To answer these questions, Tab. 12 conducts additional ablation on specific metadata components. Results show that geolocation gives the most significant improvement, followed by area and time. Interestingly, geographic coordinates perform better than Cartesian coordinates despite their distortion in high-latitude regions. Using the area corresponding to the true surface coverage (e.g., cropping and resizing make the true surface coverage smaller) is necessary, without which the performance begins to drop. Using the day of the year and absolute days above one-year-period perform similarly. We use the latter such that it’s convenient to extend to long time series in the future.
@@ B.5 Pretraining details
We pretrain Copernicus-FM on the joint 220K-grid subset of Copernicus-Pretrain, with each grid being one sample unit containing aligned images from all eight modalities. For fast data loading, we convert the raw dataset into webdataset [^1] https://github.com/webdataset/webdataset format, with one grid cell being one minimum sample in the shards. During training, one image from each modality is sampled from one grid cell to construct the input for each iteration. For S1/2, we normalize the image values with channel-wise mean and standard deviation. For S3, we multiply each channel with its corresponding scale factor [^2] https://developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S3_OLCI. For S5P, we use the raw values, and replace NaN pixels with zero. For DEM, we divide the pixel values by 10000. We apply simple data augmentations to each modality, including random resized cropping with scale [0.2,1.0][0.2,1.0][ 0.2 , 1.0 ] to its corresponding input size and random horizontal flipping. Each image comes with its metadata, including geolocation (central coordinates in lon/lat), patch area (calculated from GSD and patch size in km2), and time (number of days since a reference date 1970-01-01). The geolocation and patch area are adapted dynamically based on the cropping parameters in data augmentation. Note that despite this adaptation, due to geographical projection the patch area doesn’t strictly reflect the surface area, but is accurate enough for our pretraining purpose. While S1/2/3 images have exact acquisition dates, S5P images are monthly mean and DEM doesn’t have a specific acquisition date. Therefore, we use the first day of the month for one S5P image, and the first day of the year 2015 for all DEM images.
@@ C.2 Benchmark curation
  • Cloud-S3: This is a cloud segmentation dataset with raw images from Sentinel-3 OLCI and labels from the IdePix [66] classification algorithm. We first download a few large cloudy S3 tiles (about 4800×4004800$\times$ 4004800 × 400 pixels) distributed across the globe, and then apply the IdePix algorithm using the ESA SNAP toolbox to get multi-class cloud masks. After that, we manually check the quality of the generated masks, filter out low-quality tiles, and get seven big tiles with high-quality cloud labels. Next, we remap the label IDs, georeference the tiles to GeoTIFFs, and use GDAL to crop the large tiles into small patches with size $\times 256$ 256 pixels. We remove boundary patches filled with NaN pixels, and split all high-quality patches into 1197/399/399 train/val/test subsets. The class names for the cloud masks are: invalid, clear, cloud-sure, cloud-ambiguous, cloud-shadow, and snow-ice, of which “invalid” should be ignored during training. Apart from the multi-class labels, for each image we also have one binary cloud mask. Therefore, the Cloud-S3 dataset can serve as both a multi-class and a binary segmentation dataset.
@@
  • LC100Cls-S3 and LC100Seg-S3: These two datasets are based on Sentinel-3 OLCI images and CGLS-LC100 [13] land cover maps. CGLS-LC100 is a product in the Copernicus Global Land Service (CGLS) portfolio and delivers a global 23-class land cover map at 100 m spatial resolution. We pick the map product for 2019, and sample and download S3 images and LC100 labels for about 10K locations across the globe using GEE. For each location, we download a land cover map with about $\times 288$ 288 pixels, and four seasonal S3 OLCI images each with about $\times 96$ 96 pixels (300 m resolution). Despite using bright pixel percentage to simulate cloud filtering, the resulting images still contain a large volume of clouds. To tackle this issue, we train a cloud detection model based on the previously introduced Cloud-S3 dataset and filter out model-detected cloudy images. After a final quality check, we get about 8K samples each with a land cover map and a time series of S3 images. We divide them into 5181/1727/1727 train/val/test subsets to construct the LC100Seg-S3 dataset. For LC100Cls-S3, we integrate multi-label annotations from the land cover maps for each sample, constructing a multilabel classification dataset. Note that the number of S3 time stamps for different samples may differ because of the cloud filtering process. Apart from the time series, we also pre-select one image for each sample, constructing the “static” version of LC100Cls-S3 and LC100Seg-S3, which is the default mode in Copernicus-Bench.
@@
  • Biomass-S3: This regression dataset is based on Sentinel-3 OLCI images and CCI biomass [52]. The biomass product is part of the European Space Agency’s Climate Change Initiative (CCI) program and delivers global forest above-ground biomass at 100 m spatial resolution. We pick the product for 2020, and the layer of above ground biomass (AGB, unit: tons/ha, i.e. Mg/ha) as regression ground truth, which is defined as the mass, expressed as oven-dry weight of the woody parts (stem, bark, branches and twigs) of all living trees excluding stump and roots. We sample representative regions across the globe, and download corresponding S3 images (one for each season) from GEE and biomass maps from the CCI open data portal. We crop the S3 images into patches with about $\times 96$ 96 pixels (300 m resolution), and the corresponding biomass maps into patches with about $\times 288$ 288 pixels. Similar to LC100Cls-S3 and LC100Seg-S3, the resulting S3 images contain a large volume of clouds, thus we use again the cloud detection model to filter out cloudy images. After a final quality check, we acquire 5K samples each with a biomass map and a time series of S3 images. We divide them into 3000/1000/1000 train/val/test subsets to construct the Biomass-S3 dataset. Note that the number of S3 time stamps for different samples may differ because of the cloud filtering process. Apart from the time series, we also pre-select one image for each sample, constructing the “static” version of Biomass-S3, which is also the default mode in Copernicus-Bench.
@@
  • AQ-NO2-S5P and AQ-O3-S5P: These two regression datasets are based on Sentinel-5P NO2 and O3 images and EEA air quality data products [33]. The European Environment Agency (EEA) air quality product provides values for the human health related indicators of air pollutants at 1 km2 grid covering the whole Europe, combining monitoring air quality data in a “regression-interpolation-merging-mapping” methodology and the observational values of the air quality monitoring stations used in the interpolation. We pick the products in 2021 for NO2 (annual average concentration) and Ozone (O3, 93.2 percentile of maximum daily 8-hour means, SOMO35) as regression ground truth. We sample and download S5P NO2 (“tropospheric_NO2_column_number_density”) and O3 (“O3_column_number_density”) images from GEE, and EEA NO2 and O3 maps from EEA datahub [^3] https://www.eea.europa.eu/en/datahub. We use a sample patch size of about $\times 56$ 56 pixels for both S5P and EEA. For S5P, we download two versions: 1) annual mean, and 2) seasonal mean for each season. After filtering out NaN patches and a final quality check, we get 1480/493/494 train/val/test samples for both NO2 and O3, each with an “annual” mode of 1 S5P image and a “seasonal” mode of 4 S5P images. “Annual” is the default mode in Copernicus-Bench.
@@ C.3 Benchmark characteristics
Figs. 19, 20, 21, 22, 23, 24, 25, 26, 27 and 28 visualize some examples for each dataset in Copernicus-Bench.
@@ C.4 Benchmark implementation
We use simplified data augmentations for training sets: horizontal and vertical flipping for classification tasks, and 90$\degree$-rotation, horizontal and vertical flipping for segmentation and regression tasks. No augmentation is used for validation and testing sets. Data normalization is performed on the input according to the pretrained model’s preference. For most cases, normalization is performed by subtracting the channel-wise mean and dividing by standard deviation based on the pretrained-model-preferred statistics. If there is no preference, we recommend using the statistics calculated from the training set of each dataset as a standard. For regression tasks, we do mean/std (of the training set) normalization also on the targets to stabilize training. The predicted output is later converted back to the original scale to compute evaluation metrics.
@@ D.1 Climate prediction visualization
Figs. 30 and 32 visualize the prediction results on 10-year mean/std of the six climate parameters, comparing using the geocoordinates or a combination of coordinates and Copernicus-FM-generated grid embeddings as input data. Figs. 31 and 33 further plot the prediction error (Target-Prediction) of using only coordinates, coordinates and embeddings, and only embeddings. The figures show that using raw coordinates captures the general distribution of the climate parameters but tends to be over-smooth, while introducing EO-generated grid embeddings can capture finer details and extremes.
@@
D.2 Copernicus embedding dataset
Originally, 10K 0.25$\times$0.25 grids (with all 8 modalities) are sampled from the Copernicus-Pretrain dataset and encoded using the Copernicus-FM model to get image embeddings for each modality. The embeddings are averaged over different modalities to get one embedding vector for each grid, and later used for the climate prediction tasks to investigate the potential of bridging EO and climate. As a follow-up, we extend the embeddings to the whole globe using the full Copernicus-Pretrain dataset, constructing a ”global embedding map” at 0.25 with shape 721x1440x768 (filling empty ocean grids with 0). We term this embedding dataset Copernicus-Embed-025deg, which can be seen as a semantic map that integrates various sources of satellite observations at an extremely high compression ratio. This dataset makes it very convenient to link Earth’s surface to the atmosphere (e.g., as improved static variables adding to ERA5), unlocking new possibilities in the development of weather/climate foundation models. Fig. 34 visualizes the Copernicus-Embed-025deg dataset with top-3 principal components as RGB channels (empty ocean grids as white background).
Figure 34: Visualization of the Copernicus-Embed-025deg dataset as a global embedding map (PCA to 3-dim).